OpenStudy (anonymous):

-5x+ 8 <= x-4 <= as in less than or equal to

OpenStudy (diyadiya):

\[-5x+8\le x-4\]

OpenStudy (anonymous):

Do you want the solution ?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

$$x \ge 2$$

OpenStudy (anonymous):

i got x ≥ -1

OpenStudy (anonymous):

how did you get that answer???

OpenStudy (diyadiya):

\[-5x+8 \le x-4\] \[-5x-x \le -4 -8\] \[-6x \le -12\] \[x \ge 2\]